# Classification of abelian Schur groups I

### v1

@@ 2.4. Minimal $S$-rings
Two permutation groups $K_{1}$ and $K_{2}$ on a set $\Omega$ are called $2$-equivalent if $\operatorname{Orb}(K_{1},\Omega^{2})=\operatorname{Orb}(K_{2},\Omega^{2})$ (here we assume that $K_{1}$ and $K_{2}$ act on $\Omega^{2}$ componentwise). In this case, we write $K_{1}\approx_{2}K_{2}$. The relation $\approx_{2}$ is an equivalence relation on the set of all subgroups of $\operatorname{Sym}(\Omega)$. Every $\approx_{2}$-equivalence class has the largest element with respect to inclusion. Given $K\leq\operatorname{Sym}(\Omega)$, this element from the class containing $K$ is called the $2$-closure of $K$ and denoted by $K^{(2)}$. It should be noted that an $\approx_{2}$-equivalence class may have several minimal with respect to inclusion elements (see Example $1$).
@@
The set of all minimal with respect to inclusion permutation groups $K^{\prime}\leq\operatorname{Sym}(G)$ such that $K^{\prime}\geq G_{r}$ and $V(K^{\prime},G)=\mathcal{A}$ is denoted by $\mathcal{K}^{\min}(\mathcal{A})$.
@@
The next lemma and remark after it easily follow from computer calculations using [14].
@@
Let $|G|=4$. Then $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)\cong S_{4}$ and $\mathcal{T}_{G}=V(\operatorname{Sym}(G),G)=V(\operatorname{Alt}(G),G)$. This implies that $\mathcal{T}_{G}$ is not $2$-minimal. The group $\operatorname{Alt}(G)\cong A_{4}$ of order $12$ is a unique proper subgroup of $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)$ which is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{G})$ and hence $\operatorname{Alt}(G)\in\mathcal{K}^{\min}(\mathcal{T}_{G})$.
@@ 3.1. Tensor product
Let $\mathcal{A}$ be an $S$-ring over a group $G$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups such that $G=G_{1}\times G_{2}$. The $S$-ring $\mathcal{A}$ is defined to be a tensor product of $\mathcal{A}_{G_{1}}$ and $\mathcal{A}_{G_{2}}$ if
$\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{G_{1}})\otimes\mathcal{S}(\mathcal{A}_{G_{2}})=\{X_{1}\times X_{2}:~X_{1}\in\mathcal{S}(\mathcal{A}_{G_{1}}),~X_{2}\in\mathcal{S}(\mathcal{A}_{G_{2}})\}.$
In this case, we write $\mathcal{A}=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$. The tensor product is called nontrivial if $\{e\}<G_{1}<G$ and $\{e\}<G_{2}<G$ and trivial otherwise.
@@
It is easy to check that
(3) $\operatorname{Aut}(\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}})=\operatorname{Aut}(\mathcal{A}_{G_{1}})\otimes\operatorname{Aut}(\mathcal{A}_{G_{2}}).$
@@
A tensor product of two $S$-rings is schurian if and only if each of the operands so is.
@@
[10, Lemma 2.3] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G=G_{1}\times G_{2}$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups. Then
@@
Since $V(K,G)=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, we conclude that $\mathcal{A}_{G_{1}}=V(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=V(K_{2},G_{2})$. By the condition of the lemma, $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ which yields that $K_{i}=K^{G_{i}}$. Thus, $K=K^{G_{1}}\times K^{G_{2}}$ as required. ∎
@@ 3.2. Generalized wreath product
Let $S=U/L$ be an $\mathcal{A}$-section of $G$. The $S$-ring $\mathcal{A}$ is called the $S$-wreath product or generalized wreath product of $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ if $L\trianglelefteq G$ and every basic set $X$ of $\mathcal{A}$ outside $U$ is a union of some $L$-cosets or, equivalently, $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. In this case, we write $\mathcal{A}=\mathcal{A}_{U}\wr_{S}\mathcal{A}_{G/L}$. The $S$-wreath product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise. The construction of a generalized wreath product of $S$-rings was introduced in [4]. If $L=U$, then the $S$-wreath product coincides with the wreath product $\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ of $\mathcal{A}_{L}$ and $\mathcal{A}_{G/L}$. The $S$-ring $\mathcal{A}$ is said to be decomposable if $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S$ and indecomposable otherwise.
@@
[6, Corollary 5.7] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$ and $S=U/L$ an $\mathcal{A}$-section. Suppose that $\mathcal{A}$ is the $S$-wreath product and $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian. Then $\mathcal{A}$ is schurian if and only if there exist two groups $K_{1}$ and $K_{0}$ such that
@@
[21, Corollary 10.3] Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian if $\mathcal{A}_{S}$ is $2$-minimal. In particular, $\mathcal{A}$ is schurian if $U=L$.
The lemma below for cyclic groups was proved in [9, Theorem 7.5]. In fact, the proof for abelian groups is the same. To make the text self-contained, we provide a proof here.
@@ 3.3. Star product
In this case, we write $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. The construction of a star product of $S$-rings was introduced in [12] and extended to association schemes under the name “crested product” in [1]. A star product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise.
@@ 4. Duality
For any $S$-ring $\mathcal{A}$ over the group $G$, one can define the dual $S$-ring $\widehat{\mathcal{A}}$ over $\widehat{G}$ as follows: two irreducible characters of $G$ belong to the same basic set of $\widehat{\mathcal{A}}$ if they have the same value on each basic set of $\mathcal{A}$ (for the exact definition, we refer the readers to [7, 31]). One can verify that
@@
and the $S$-ring dual to $\widehat{\mathcal{A}}$ is equal to $\mathcal{A}$. The following lemma collects some facts on the dual $S$-ring taken from [8, Section 2.3].
@@
    $\mathcal{A}$ is the $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ if and only if $\widehat{\mathcal{A}}$ is the $L^{\bot}/U^{\bot}$-wreath product;
@@ 5. $S$-rings over $H\times C_{p}$
[17, Lemma 6.3] If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, the Hall $p^{\prime}$-subgroup of $H_{1}P_{1}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
@@
[22, Proposition 15] If $\mathcal{A}_{H_{1}P_{1}/H_{1}}\cong\mathbb{Z}C_{p}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
@@
    $H_{1}P_{1}=G$, $P_{1}<G$, and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$;
@@
From [10, Lemma 6.2] it follows that Statement $(2)$ of the lemma holds or $\mathcal{A}$ is the $(H_{1}P_{1})/P_{1}$-wreath product with $P_{1}<G$. If the above wreath product is nontrivial, then Statement $(3)$ of the lemma holds, whereas if it is trivial, i.e. if $H_{1}P_{1}=G$, then $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}=G_{p^{\prime}}=H$. So $\mathcal{A}=\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$ by Lemma 5.1 and Statement $(1)$ of the lemma holds. ∎
@@ 6. $S$-rings over cyclic groups
In this section, we provide several results on $S$-rings over cyclic groups required for the proofs of the main results. An $S$-ring over a cyclic group is said to be circulant.
@@ 6.1. General results
[32, Theorem 25.4] If $G$ is an abelian group of composite order having a cyclic Sylow subgroup, then $\mathcal{T}_{G}$ is the only primitive $S$-ring over $G$.
The description of $S$-rings over a cyclic group was obtained in [18, 19]. Below, we provide this description in a convenient for us form taken from [6, Theorem 4.1]. If $\mathcal{A}$ is an $S$-ring over a cyclic group $G$, then put $\operatorname{rad}(\mathcal{A})=\operatorname{rad}(X)$, where $X$ is a basic set of $\mathcal{A}$ containing a generator of $G$. Due to Lemma 2.3, all basic sets of $\mathcal{A}$ containing a generator of $G$ are rationally conjugate and hence have the same radical. Therefore $\operatorname{rad}(\mathcal{A})$ does not depend on the choice of $X$.
@@
    $|\operatorname{rad}(\mathcal{A})|=1$ if and only if $\mathcal{A}$ is a tensor product of a normal cyclotomic $S$-ring with trivial radical and trivial $S$-rings;
@@
[6, Theorem 4.2(1)] Every normal $S$-ring over a cyclic group is cyclotomic and, in particular, schurian.
@@
Suppose that $\mathcal{A}_{S}=\mathcal{T}_{S}$. Then $\mathcal{A}_{S}=\operatorname{Cyc}(\operatorname{Aut}(S),S)=V(\operatorname{Hol}(S),S)$. Since $\operatorname{Aut}(\mathcal{A}_{U})$ and $\mathcal{A}_{G/L}$ are schurian, [6, Theorem 8.1(2)] implies that there exist groups $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that
@@ 6.2. $S$-rings over cyclic $p$-groups
The lemma below can be found, e.g., in [23, Lemma 2.4].
@@
The first statement of the lemma below immediately follows from [5, Lemma 5.1], whereas the second one follows from Lemma 6.2 and the first one.
@@
Let $p$ be an odd prime and $G$ a cyclic $p$-group.
@@
    If $K\leq\operatorname{Aut}(G)$ and $|\operatorname{rad}(X)|=1$ for some $X\in\operatorname{Orb}(K,G)$ such that $\langle X\rangle=G$, then $|Y|=|K|\leq p-1$ and all the elements of $Y$ lie in pairwise distinct cosets by every proper subgroup of $\langle Y\rangle$ for every $Y\in\operatorname{Orb}(K,G)\setminus\{\{e\}\}$.
@@
    If $\mathcal{A}$ is a nontrivial $S$-ring over $G$ such that $|\operatorname{rad}(\mathcal{A})|=1$, then $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ with $|K|$ dividing p1p-1 and $|X|=|K|$ for every $X\in\mathcal{S}(\mathcal{A})\setminus\{\{e\}\}$.
@@
Since $G$ is a cyclic $p$-group, there exists the least $\mathcal{A}$-subgroup $L$. Let $U$ be the $\mathcal{A}$-subgroup generated by all basic set of $\mathcal{A}$ with trivial radical. The statement of the lemma immediately follows from the definitions of $L$ and $U$. ∎
@@
Let $\mathcal{A}$ be an $S$-ring over a cyclic $p$-group $G$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X$ is not regular. Then $X=U\setminus L$ for some $\mathcal{A}$-subgroups $U>L$ and $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}=\mathcal{A}_{U}\wr\mathcal{A}_{G/U}$.

### v3

@@ 2.4. Minimal $S$-rings
Two permutation groups $K_{1}$ and $K_{2}$ on a set $\Omega$ are called $2$-equivalent if $\operatorname{Orb}(K_{1},\Omega^{2})=\operatorname{Orb}(K_{2},\Omega^{2})$ (here we assume that $K_{1}$ and $K_{2}$ act on $\Omega^{2}$ componentwise). In this case, we write $K_{1}\approx_{2}K_{2}$. The relation $\approx_{2}$ is an equivalence relation on the set of all subgroups of $\operatorname{Sym}(\Omega)$. Every $\approx_{2}$-equivalence class has the unique maximal element and may have more than one minimal elements with respect to inclusion (for the latter, see Example $1$). Given $K\leq\operatorname{Sym}(\Omega)$, the unique maximal element from the class containing $K$ is called the $2$-closure of $K$ and denoted by $K^{(2)}$.
@@
The set of all minimal with respect to inclusion permutation groups $K^{\prime}\leq\operatorname{Sym}(G)$ such that $K^{\prime}\geq G_{r}$ and $V(K^{\prime},G)=\mathcal{A}$ is denoted by $\mathcal{K}^{\min}(\mathcal{A})$. Note that $\mathcal{A}$ is $2$-minimal if and only if $\operatorname{Aut}(\mathcal{A})$ is $2$-isolated in the sense of [15].
@@
The next lemma and remark after it easily follow from computer calculations using the package [16].
@@
Let $|G|=4$. Then $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)\cong\operatorname{Sym}(4)$ and $\mathcal{T}_{G}=V(\operatorname{Sym}(G),G)=V(\operatorname{Alt}(G),G)$. This implies that $\mathcal{T}_{G}$ is not $2$-minimal. The group $\operatorname{Alt}(G)\cong\operatorname{Alt}(4)$ of order $12$ is a unique proper subgroup of $\operatorname{Aut}(\mathcal{T}_{G})=\operatorname{Sym}(G)$ which is $2$-equivalent to $\operatorname{Aut}(\mathcal{T}_{G})$ and hence $\mathcal{K}^{\min}(\mathcal{T}_{G})=\{\operatorname{Alt}(G)\}$.
@@ 3.1. Tensor product
Let $\mathcal{A}$ be an $S$-ring over a group $G$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups such that $G=G_{1}\times G_{2}$. The $S$-ring $\mathcal{A}$ is defined to be a tensor product of $\mathcal{A}_{1}=\mathcal{A}_{G_{1}}$ and $\mathcal{A}_{2}=\mathcal{A}_{G_{2}}$ if
$\mathcal{S}(\mathcal{A})=\mathcal{S}(\mathcal{A}_{1})\otimes\mathcal{S}(\mathcal{A}_{2})=\{X_{1}\times X_{2}:~X_{1}\in\mathcal{S}(\mathcal{A}_{1}),~X_{2}\in\mathcal{S}(\mathcal{A}_{2})\}.$
In this case, we write $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$. The tensor product is called nontrivial if $\{e\}<G_{1}<G$ and $\{e\}<G_{2}<G$, and trivial otherwise. On the other hand, if we are given $S$-rings $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ over $G_{1}$ and $G_{2}$, respectively, then the above partition defines the $S$-ring $\mathcal{A}=\mathcal{A}_{1}\otimes\mathcal{A}_{2}$ over $G=G_{1}\times G_{2}$.
@@
One can check that
(3) $\operatorname{Aut}(\mathcal{A}_{1}\otimes\mathcal{A}_{2})=\operatorname{Aut}(\mathcal{A}_{1})\otimes\operatorname{Aut}(\mathcal{A}_{2}).$
@@
A tensor product of two $S$-rings is schurian if and only if each of them is schurian.
@@
[11, Lemma 2.3] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G=G_{1}\times G_{2}$. Suppose that $G_{1}$ and $G_{2}$ are $\mathcal{A}$-subgroups. Then
@@
Since $V(K^{G_{1}},G_{1})=\mathcal{A}_{G_{1}}$, $V(K^{G_{2}},G_{2})=\mathcal{A}_{G_{2}}$, and $V(K,G)=\mathcal{A}_{G_{1}}\otimes\mathcal{A}_{G_{2}}$, we conclude that $\mathcal{A}_{G_{1}}=V(K_{1},G_{1})$ and $\mathcal{A}_{G_{2}}=V(K_{2},G_{2})$. By the condition of the lemma, $K^{G_{i}}\in\mathcal{K}^{\min}(\mathcal{A}_{G_{i}})$ which yields that $K_{i}=K^{G_{i}}$. Thus, $K=K^{G_{1}}\times K^{G_{2}}$ as required. ∎
@@ 3.2. Generalized wreath product
Let $S=U/L$ be an $\mathcal{A}$-section of $G$. The $S$-ring $\mathcal{A}$ is called the $S$-wreath product or generalized wreath product of $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ if $L\trianglelefteq G$ and every basic set $X$ of $\mathcal{A}$ outside $U$ is a union of some $L$-cosets or, equivalently, $L\leq\operatorname{rad}(X)$ for every $X\in\mathcal{S}(\mathcal{A})_{G\setminus U}$. In this case, we write $\mathcal{A}=\mathcal{A}_{U}\wr_{S}\mathcal{A}_{G/L}$. The $S$-wreath product is said to be nontrivial if $L\neq\{e\}$ and $U\neq G$, and trivial otherwise. The construction of a generalized wreath product of $S$-rings was introduced in paper [5]. If $L=U$, then the $S$-wreath product coincides with the wreath product $\mathcal{A}_{L}\wr\mathcal{A}_{G/L}$ of $\mathcal{A}_{L}$ and $\mathcal{A}_{G/L}$. The $S$-ring $\mathcal{A}$ is said to be decomposable if $\mathcal{A}$ is a nontrivial $S$-wreath product for some $\mathcal{A}$-section $S$ and indecomposable otherwise.
@@
[7, Corollary 5.7] Let $\mathcal{A}$ be an $S$-ring over an abelian group $G$ and $S=U/L$ an $\mathcal{A}$-section. Suppose that $\mathcal{A}$ is the $S$-wreath product and $\mathcal{A}_{U}$ and $\mathcal{A}_{G/L}$ are schurian. Then $\mathcal{A}$ is schurian if and only if there exist two groups $K_{1}\leq\operatorname{Sym}(U)$ and $K_{0}\leq\operatorname{Sym}(G/L)$ such that
@@
[23, Corollary 10.3] Under the hypothesis of Lemma 3.5, the $S$-ring $\mathcal{A}$ is schurian if $\mathcal{A}_{S}$ is $2$-minimal. In particular, $\mathcal{A}$ is schurian if $U=L$.
The lemma below for cyclic groups is [10, Theorem 7.5]. In fact, the proof for abelian groups is the same. To make the text self-contained, we provide a proof here.
@@ 3.3. Star product
In this case, we write $\mathcal{A}=\mathcal{A}_{L}\star\mathcal{A}_{U}$. The construction of a star product of $S$-rings was introduced in paper [13] and extended to association schemes under the name “crested product” in paper [1]. A star product is called nontrivial if $L\neq\{e\}$ and $U\neq G$ and trivial otherwise.
@@ 4. Duality
For any $S$-ring $\mathcal{A}$ over the group $G$, one can define the dual $S$-ring $\widehat{\mathcal{A}}$ over $\widehat{G}$ as follows: two irreducible characters of $G$ belong to the same basic set of $\widehat{\mathcal{A}}$ if they have the same value on each basic set of $\mathcal{A}$ (for the exact definition, we refer the reader to [8, 33]). One can verify that
@@
and the $S$-ring dual to $\widehat{\mathcal{A}}$ is equal to $\mathcal{A}$. The following lemma collects some facts on the dual $S$-ring taken from [9, Section 2.3].
@@
    $\mathcal{A}$ is the $U/L$-wreath product for some $\mathcal{A}$-section $U/L$ if and only if $\widehat{\mathcal{A}}$ is the $L^{\bot}/U^{\bot}$-wreath product.
@@ 5. $S$-rings over $H\times C_{p}$
[19, Lemma 6.3] If $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}$, the Hall $p^{\prime}$-subgroup of $H_{1}P_{1}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
@@
[24, Proposition 15] If $\mathcal{A}_{(H_{1}P_{1})/H_{1}}\cong\mathbb{Z}C_{p}$, then $\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$.
@@
    $H_{1}P_{1}=G$, $P_{1}\lneq G$, and $\mathcal{A}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$;
@@
From [11, Lemma 6.2] it follows that Statement $(2)$ of the lemma holds or $\mathcal{A}$ is the $(H_{1}P_{1})/P_{1}$-wreath product with $P_{1}<G$. If the above wreath product is nontrivial, then Statement $(3)$ of the lemma holds, whereas if it is trivial, i.e. if $H_{1}P_{1}=G$, then $H_{1}\neq(H_{1}P_{1})_{p^{\prime}}=G_{p^{\prime}}=H$. So $\mathcal{A}=\mathcal{A}_{H_{1}P_{1}}=\mathcal{A}_{H_{1}}\star\mathcal{A}_{P_{1}}$ by Lemma 5.1 and Statement $(1)$ of the lemma holds. ∎
@@ 6. $S$-rings over cyclic groups
In this section, we provide several results on $S$-rings over cyclic groups, required for the proofs of the main results. An $S$-ring over a cyclic group is said to be circulant.
@@ 6.1. General results
[34, Theorem 25.4] If $G$ is an abelian group of composite order having a cyclic Sylow subgroup, then $\mathcal{T}_{G}$ is the only primitive $S$-ring over $G$.
The description of $S$-rings over a cyclic group was obtained in papers [20, 21]. Below, we provide this description in a convenient for us form taken from [7, Theorem 4.1]. If $\mathcal{A}$ is an $S$-ring over a cyclic group $G$, then put $\operatorname{rad}(\mathcal{A})=\operatorname{rad}(X)$, where $X$ is a basic set of $\mathcal{A}$ containing a generator of $G$. Due to Lemma 2.3, all basic sets of $\mathcal{A}$, containing a generator of $G$, are rationally conjugate and hence have the same radical. Therefore $\operatorname{rad}(\mathcal{A})$ does not depend on the choice of $X$.
@@
    $|\operatorname{rad}(\mathcal{A})|=1$ if and only if $\mathcal{A}$ is a tensor product (possibly, trivial) of a normal cyclotomic $S$-ring with trivial radical and trivial $S$-rings;
@@
[7, Theorem 4.2(1)] Every normal $S$-ring over a cyclic group is cyclotomic and, in particular, schurian.
@@
Suppose that $\mathcal{A}_{S}=\mathcal{T}_{S}$. Then $\mathcal{A}_{S}=\operatorname{Cyc}(\operatorname{Aut}(S),S)=V(\operatorname{Hol}(S),S)$. Since $\operatorname{Aut}(\mathcal{A}_{U})$ and $\mathcal{A}_{G/L}$ are schurian, [7, Theorem 8.1(2)] implies that there exist groups $K_{1}\leq\operatorname{Aut}(\mathcal{A}_{U})$ and $K_{0}\leq\operatorname{Aut}(\mathcal{A}_{G/L})$ such that
@@ 6.2. $S$-rings over cyclic $p$-groups
For the lemma below see, e.g., [25, Lemma 2.4].
@@
The first statement of the lemma below immediately follows from [6, Lemma 5.1], whereas the second one follows from Lemma 6.2 and the first one.
@@
Let $p$ be an odd prime, $G$ a cyclic $p$-group, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})\setminus\{\{e\}\}$.
@@
    If $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ and $|\operatorname{rad}(\mathcal{A})|=1$, then $|X|=|K|\leq p-1$ and all the elements of $X$ lie in pairwise distinct cosets by every proper subgroup of $\langle X\rangle$.
@@
    If $\mathcal{A}$ is nontrivial with trivial radical, then $\mathcal{A}=\operatorname{Cyc}(K,G)$ for some $K\leq\operatorname{Aut}(G)$ such that $|K|$ divides p1p-1 and $|X|=|K|$.
@@
Since $G$ is a cyclic $p$-group, there exists the least nontrivial $\mathcal{A}$-subgroup $L$. Let $U$ be the $\mathcal{A}$-subgroup generated by all basic set of $\mathcal{A}$ with trivial radical. Note that $U<G$ because $\mathcal{A}$ is a nontrivial generalized wreath product. Every basic set $X$ of $\mathcal{A}$ outside $U$ has a nontrivial radical and hence $\operatorname{rad}(X)\geq L$. Thus, $\mathcal{A}$ is the nontrivial $U/L$-wreath product as desired. ∎
@@
Let $\mathcal{A}$ be an $S$-ring over a cyclic $p$-group $G$ and $X\in\mathcal{S}(\mathcal{A})$. Suppose that $X$ is not regular. Then $X=U\setminus L$ for some $\mathcal{A}$-subgroups $U>L$ such that $|U:L|\geq p^{2}$ and $\mathcal{A}=\mathcal{A}_{L}\wr\mathcal{A}_{G/L}=\mathcal{A}_{U}\wr\mathcal{A}_{G/U}$.